# Toward Any-to-Any Emotion Voice Conversion using Disentangled Diffusion Framework

### v2

@@ Abstract
Emotional Voice Conversion (EVC) aims to modify the emotional expression of speech for various applications, such as human-machine interaction. Previous deep learning-based approaches using generative adversarial networks and autoencoder models have shown promise but suffer from quality degradation and limited emotion control. To address these issues, a novel diffusion-based EVC framework with disentangled loss and expressive guidance is proposed. Our method separates speaker and emotional features to maintain speech quality while enhancing emotional expressiveness. Tested on real-world and acted-out datasets, the approach achieved significant improvements in emotion classification accuracy for both in-the-wild and act-out datasets and showed reduced distortion compared to state-of-the-art models.
@@ I Introduction
Emotional Voice Conversion (EVC), which focuses on artificially modifying the emotional expression of speech signals, offers potential applications such as enhancing the naturalness of human-machine interaction [1, 2]. By increasing the controllability over complex factors, including voice quality, emotional expressiveness, speaker traits, and linguistic contents, for human speech generation, we can gradually perform one-to-one, many-to-many, and finally, any-to-any emotion voice conversion, which is applicable to convert the emotion of unseen speakers. With the effectiveness of deep learning techniques, building and evaluating the EVC model on specific target speakers or small-scaled parallel emotional speech has gained much success in previous works [3, 4, 5, 6].
@@
When addressing various EVC issues, most of the deep learning methods models lie in the two realms, which are generative adversarial networks [3, 4] and autoencoder-based models [5, 6]. GAN-based methods leverage adversarial mechanisms to learn direct mappings between data distributions of different emotional classes. Autoencoder-based approaches tackle this issue by introducing a disentanglement mechanism that separates linguistic and speaker identity features from emotional representations, enabling better control over emotion conversion. However, these works are suboptimal in converting the source emotion to the target emotion and can distort the converted voice, i.e., degrading the generated voice’s naturalness and emotion quality.
@@
The diffusion model has recently garnered significant attention for its generative capabilities of high-quality samples in multiple application scenarios [7, 8, 9]. A most recent emotion conversion work, EmoConv-Diff [10] demonstrated a statistically significant improvement in intensity controllability compared to their previous method [11] while achieving great quality. However, the controllability of valence is still challenging, and it is unclear whether the speaker and emotion information are learned and extracted from the corresponding entangled representation. To tackle the problems, we design our framework with disentanglement in two separate aspects to generate better emotional characteristics while sustaining high speech quality comparable to original source utterances. First, we incorporate the disentangle loss into the diffusion EVC model training process. Second, with the reversed diffusion process, we further design the expressive guidance mechanism to enhance the expressiveness of the target emotion and mitigate the distortion of the source speaker trait.
@@
In this work, different from working in the controlled scenarios such as [3, 4, 5, 6], we aim to solve any-to-any emotional voice conversion by proposing a novel diffusion framework design with disentangled loss and expressive guidance. To demonstrate the effectiveness, we train our model on un-parallel real-world speech emotion data (MSP-Podcast v1.10 [12]) and rigorously evaluate our method on two different corpora, including the real-world scenario (MSP-Podcast v1.11) and act-out scenario (ESD [1]), with the unseen environment, speakers and linguistic contents. Both objective and subjective evaluations show that our proposed method effectively converts emotion in speech for in-the-wild data with 21% ECA improvement on MSP and 32% ECA improvement on ESD. Moreover, with objective metrics of 3.718 in UTMOS, 3.600 in SIG, and subjective metrics of 4.024 in nMOS, and 3.159 in sMOS, our experiment results demonstrate that our proposed any-to-any EVC can generate emotional speech comparable to state-of-the-art EVC with significantly less distortion.
@@ II-A Proposed Method
Figure 1 shows the overall framework of our proposed method, which comprises a set of encoders modeling each component, a diffusion-based decoder, a disentanglement mechanism, and a guidance method. We use pre-trained HiFiGAN vocoder [13] to convert the mel spectrogram generated by the proposed framework back to the time domain signal.
@@ II-A1 Encoders
Three pre-trained encoders are used to capture linguistic content representations $c$, speaker identity $spk$, and emotional expression $emo$.
@@ II-A4 Disentangled Loss
where $\lambda_{MI}$ and $\lambda_{guide}$ are hyparameters to control the importance of respective loss.
@@ III-A2 Evaluation Setup
First, we perform any-to-any emotion voice conversion that includes all of the transformations between angry, happy, sad, and neutral, besides transforming from emotional speech to neutral one. The experiments are conducted on both in-the-wild datasets (MSP-Podcast) and high-quality act-out datasets (ESD). We compared methods under different training schemes, i.e., using only $\mathcal{L}_{diff}$ or using $\mathcal{L}_{Total}$ in equation 2. We then apply the proposed expressive guidance method on the model trained with $\mathcal{L}_{Total}$. We compared $EG_{spk}$, $EG_{emo}$, and $EG_{spk,emo}$ with different settings of negative condition that replace the representation of positive condition corresponding to either speaker identity $spk$, emotion information $emo$ or both. Second, We compared our proposed method with $EG_{spk,emo}$ against five different models, i.e. CycleGAN-EVC [3], StarGAN-EVC [4], Seq2Seq-EVC [5], Emovox [6] and Prosody2Vec [23] following conversion samples based on ESD presented in Prosody2Vec [^1] https://leyuanqu.github.io/Prosody2Vec/. Unlike ours, models in comparison are trained on act-out ESD datasets, while only Prosody2Vec utilized both predominant and in-the-wild datasets. The audio samples are available on our demo page [^2] https://henrychou36.github.io/Any-to-Any-EVC/.
@@ III-A3 Evaluation Metric
For both experiments, we incorporate non-intrusive objective evaluation, i.e., UTMOS [24] for natureliness, DNSMOS [25] for speech quality (SIG) and overall signal quality (OVRL). Both methods are designed to predict the mean opinion score (MOS) of subjective listening tests. To access speaker similarity, speaker embedding cosine similarity (SECS) between extracted embeddings of source and generated speech based on Resemblyzer [15] is used. For emotion classification accuracy (ECA), we utilized a speech emotion recognition (SER) model pre-trained on both MSP-Podcast and ESD based on emotion embedding from [16]. For the second experiment, in addition to objective evaluation, we conducted a subjective assessment with 14 subjects evaluating 63 converted or target utterances using a 5-point scale ranging from 1 to 5 to assess speech quality, naturalness, and emotion similarity between synthesis speech and the target utterances. We report the mean opinion scores with a 95% confidence interval for speech quality (MOS), naturalness (nMOS), and emotion similarity (sMOS). The subjects are also required to label the primary emotion for ECA. The evaluation of the first and second experiments are presented in Table I and Table II, separately.
@@
Emovox many-to-mnay 2.381 0.698 3.234 2.930 0.333 2.683$\pm$0.202 2.191$\pm$0.158 2.508$\pm$0.235 0.325
@@ IV Conclusion and future work
It is important to develop ”any-to-any” EVC for real-world applications. In this work, we propose an any-to-any emotion voice conversion that combines a disentanglement mechanism and expressive guidance and provides thorough evaluation with both objective and subjective tests over both in-the-wild and act-out datasets. We show that our proposed framework can effectively convert speech into different emotions while having high speech and audio quality. Moreover, compared to other EVC frameworks in the control environment, our method generates distortionless emotional speech. We also address the challenge of controllability over valence, which will be further evaluated and developed for our future works.

### v3

@@ Abstract
Emotional Voice Conversion (EVC) aims to modify the emotional expression of speech for various applications, such as human-machine interaction. Previous deep learning-based approaches using generative adversarial networks and autoencoder models have shown promise but suffer from quality degradation and limited emotion control. To address these issues, a novel diffusion-based EVC framework with disentangled loss and expressive guidance is proposed. Our method separates speaker and emotional features to maintain speech quality while improving emotional expressiveness. Tested on real-world and acted-out datasets, the approach achieved significant improvements in emotion classification accuracy for both in-the-wild and act-out datasets and showed reduced distortion compared to state-of-the-art models.
@@ I Introduction
Emotional Voice Conversion (EVC), which focuses on artificially modifying the emotional expression of speech signals, offers potential applications such as improving the naturalness of human-machine interaction [1, 2]. By increasing the controllability over complex factors, including voice quality, emotional expressiveness, speaker traits, and linguistic contents, for human speech generation, we can gradually perform one-to-one, many-to-many, and finally, any-to-any emotion voice conversion, which is applicable to convert the emotion of unseen speakers. With the effectiveness of deep learning techniques, building and evaluating the EVC model with specific target speakers or small-scale parallel emotional speech has gained much success in previous works [3, 4, 5, 6].
@@
When addressing various EVC issues, most deep learning methods models lie in the two realms, which are generative adversarial networks [3, 4] and autoencoder-based models [5, 6]. GAN-based methods leverage adversarial mechanisms to learn direct mappings between data distributions of different emotional classes. Autoencoder-based approaches tackle this issue by introducing a disentanglement mechanism that separates linguistic and speaker identity features from emotional representations, enabling better control over emotion conversion. However, these works are suboptimal in converting the source emotion to the target emotion and can distort the converted voice, i.e., degrading the generated voice’s naturalness and emotion quality.
@@
The diffusion model has recently gained significant attention for its generative capabilities of high-quality samples in multiple application scenarios [7, 8, 9]. A most recent emotion conversion work, EmoConv-Diff [10] demonstrated a statistically significant improvement in intensity controllability compared to their previous method [11] while achieving great quality. However, the controllability of valence is still challenging, and it is unclear whether the speaker and emotion information are learned and extracted from the corresponding entangled representation. To tackle the problems, we design our framework with disentanglement in two separate aspects to generate better emotional characteristics while sustaining high speech quality comparable to source utterances. First, we incorporate the disentangle loss into the diffusion EVC model training process. Second, with the reversed diffusion process, we further design the expressive guidance mechanism to enhance the expressiveness of the target emotion and mitigate the distortion of the source speaker trait.
@@
In this work, different from working in controlled scenarios such as [3, 4, 5, 6], we aim to solve any-to-any emotional voice conversion by proposing a novel diffusion framework design with disentangled loss and expressive guidance. To demonstrate the effectiveness, we train our model on nonparallel real-world speech emotion data (MSP-Podcast v1.10 [12]) and rigorously evaluate our method on two different corpora, including the real-world scenario (MSP-Podcast v1.11) and act-out scenario (ESD [13]), with the unseen environment, speakers and linguistic contents. Both objective and subjective evaluations show that our proposed method effectively converts emotion in speech for in-the-wild data with 21% ECA improvement on MSP and 32% ECA improvement on ESD datasets. Furthermore, with objective metrics of 3.718 in UTMOS, 3.600 in SIG, and subjective metrics of 4.024 in nMOS, and 3.159 in sMOS, our experimental results demonstrate that our proposed any-to-any EVC can generate emotional speech comparable to state-of-the-art EVC with significantly less distortion.
@@ II-A Proposed Method
Figure 1 shows the overall framework of our proposed method, which comprises a set of encoders modeling each component, a diffusion-based decoder, a disentanglement mechanism, and a guidance method. We utilize pre-trained HiFiGAN vocoder [14] to convert the mel spectrogram generated by the proposed framework back to the time domain signal.
@@ II-A1 Encoders
Three pre-trained encoders capture linguistic content representations $c$, speaker identity $spk$, and emotional expression $emo$.
@@ II-A4 Disentangled Loss
where $\lambda_{MI}$ and $\lambda_{style}$ are hyparameters to control the importance of respective loss.
@@ III-A2 Evaluation Setup
First, we performed any-to-any emotion voice conversion that includes all of the transformations between angry, happy, sad, and neutral, besides transforming from emotional speech to neutral one. The experiments are conducted on both in-the-wild datasets (MSP-Podcast) and high-quality act-out datasets (ESD). We compared methods under different training schemes, i.e., using only $\mathcal{L}_{diff}$ or using $\mathcal{L}_{Total}$ in equation 2. We then apply the proposed expressive guidance method on the model trained with $\mathcal{L}_{Total}$. We compared $EG_{spk}$, $EG_{emo}$, and $EG_{spk,emo}$ with different settings of negative condition that replace the representation of positive condition corresponding to either speaker identity $spk$, emotion information $emo$ or both. Second, We compared our proposed method with $EG_{spk,emo}$ against five different models, i.e. CycleGAN-EVC [3], StarGAN-EVC [4], Seq2Seq-EVC [5], Emovox [6] and Prosody2Vec [23] following conversion samples based on ESD presented in Prosody2Vec [^1] https://leyuanqu.github.io/Prosody2Vec/. Unlike ours, models in comparison are trained on act-out ESD datasets, while only Prosody2Vec utilized both predominant and in-the-wild datasets. The audio samples are available on our demo page [^2] https://henrychou36.github.io/Any-to-Any-EVC/.
@@ III-A3 Evaluation Metric
For both experiments, we incorporate non-intrusive objective evaluation, i.e., UTMOS [24] for naturelness, DNSMOS [25] for speech quality (SIG) and overall signal quality (OVRL). Both methods are designed to predict the mean opinion score (MOS) of subjective listening tests. To access speaker similarity, speaker embedding cosine similarity (SECS) between extracted embeddings of source and generated speech based on Resemblyzer [16] is used. For emotion classification accuracy (ECA), we utilized a speech emotion recognition (SER) model pre-trained on both MSP-Podcast and ESD based on emotion embedding from [17]. For the second experiment, in addition to objective evaluation, we conducted a subjective assessment with 14 subjects evaluating 63 converted or target utterances using a 5-point scale ranging from 1 to 5 to assess speech quality, naturalness, and emotion similarity between synthesis speech and the target utterances. We report the mean opinion scores with a 95% confidence interval for speech quality (MOS), naturalness (nMOS), and emotion similarity (sMOS). The subjects are also required to label the primary emotion for ECA. The evaluation of the first and second experiments are presented in Table I and Table II, separately.
@@
Emovox many-to-many 2.381 0.698 3.234 2.930 0.333 2.683$\pm$0.202 2.191$\pm$0.158 2.508$\pm$0.235 0.325
@@ IV Conclusion and future work
It is important to develop ”any-to-any” EVC for real-world applications. In this work, we propose an any-to-any emotion voice conversion that combines a disentanglement mechanism and expressive guidance and provides thorough evaluation with both objective and subjective tests over both in-the-wild and act-out datasets. We show that our proposed framework can effectively convert speech into different emotions while having high speech and audio quality. Moreover, our method generates distortionless emotional speech compared to other EVC frameworks in the control environment. We also address the challenge of controllability over valence, which will be further evaluated and developed for our future works.